2.6.17. The probability density function of the random variable X is given by r2 21 0<x-1, 6x-2r2-3 (x, 3)2 0 otherwise.
Answer #1

E(X) = int x f(x) dx

-2 - (6r - 22 -3) 2. (2-3 0

2

Т (62.-22.2-3

2. (2.-3)2

hence

E(X) = 1/8 + 1 + 3/8 = 3/2